OpenStudy (anonymous):

find the range of a function.....a)f(x)=10-x^2 b) (x^2)/(1-x^2)=f(x)

OpenStudy (anonymous):

a) is a parabola that opens downward with vertex (0, 10), so the range is from negative infinity up to the vertex\[(-\infty, 10]\]

OpenStudy (anonymous):

what is a pabola?

OpenStudy (anonymous):

b) this rational function has a horizontal asymptote y=-1, so the range is all real numbers except -1:\[(-\infty,-1) \cup (-1, \infty)\]

OpenStudy (anonymous):

parabola?*

OpenStudy (anonymous):

when you have polynomial function with x^2 it has the shape called a parabola; i will attach a sketch

OpenStudy (anonymous):

to find range you really have to know something about the graph of the function

OpenStudy (anonymous):

but isnt 10-x^2 a binomial?...im confused sry

OpenStudy (anonymous):

yes, why does that concern you?

OpenStudy (anonymous):

b.c you said its a polynomial lol

OpenStudy (anonymous):

parabolas come from functions that look like\[y=ax^2+bx+c\]where a can't equal zero, but b or c can equal zero; in your case b=0

OpenStudy (anonymous):

yes, why does that concern you?

OpenStudy (anonymous):

ohhh ok i get it thanks soo much!! =]